# Market Reaction to News Flows in Supply Chain Networks

### v3

@@ Abstract
This study examines how positive and negative news about firms affects their stock prices and, moreover, how it affects stock prices of the firms’ suppliers and clients, using a large sample of publicly listed firms around the world and another of Japanese listed firms. The level of positiveness and negativeness of each news article is determined by FinBERT, a natural language processing model fine-tuned specifically for financial information. Supply chains of firms across the world are identified mostly by financial statements, while those of Japanese firms are taken from large-scale firm-level surveys. We find that positive news increases the change rate of stock prices of firms mentioned in the news before its disclosure, most likely because of diffusion of information through unofficial information channels. Positive news also raises stock prices of the firms’ suppliers and clients before and after its disclosure, confirming propagation of market values through supply chains. In addition, we generally find a larger post-news effect on stock prices of the mentioned firms and their suppliers and clients than the pre-news effect. The positive difference between the post- and pre-news effects can be considered as the net effect of the disclosure of positive news, controlling for information diffusion through private channels. However, the post-news effect on suppliers and clients in Japan is smaller than the pre-news effect, which is the opposite result to non-domestic firms from around the world.
keywords:
Supply Chain, Natural Language Processing, Stock Market, Large -scale Data1 Introduction
@@
Firms interact with and influence each other through various types of networks. Major firm networks that are found to play a significant role in firms’ behaviors and performance in the literature include financial networks through ownership [1, 2, 3] and knowledge networks through research collaboration.[4, 5, 6, 7, 8] Another type of network recently paid attention in the academic literature and business and policy fields is supply chains formed through transactions of materials, parts, and components.[9, 10] One reason for the growing attention is that supply chains have expanded globally, linking firms in the world within a small number of steps with each other. [11]
@@
Global supply chains have rapidly expanded because interconnected firms through supply chains can greatly benefit from the network. Most notably, supply chains can empower firms by enhancing efficiency of production processes and thus optimizing firms’ economic performance.[12, 13] For instance, by procuring and manufacturing different goods in various locations based on each location’s comparative advantages and cost effectiveness, firms can streamline their operations as a network, ultimately leading to reductions in production costs and improvements in productivity.[14] Moreover, it is often evidenced that productivity spills over through supply chains from upstream suppliers to downstream client firms because of high-quality materials, parts, components, and services and from downstream to upstream firms because of learning of knowledge and technology.[15, 7, 16]
@@
Conversely, supply chains can be a channel of negative shocks that aggravate firm performance, particularly once supply chains are disrupted due to unexpected events, such as natural disasters and geopolitical instability.[17, 18, 19] Such negative shocks examined in the literature include the Great East Japan Earthquake in 2011,[20] the massive flood in Thailand in 2011,[21] the COVID-19 pandemic from 2020 to 2022,[22], and the Russia-Ukraine war that started in 2022.[23]
@@
Therefore, positive or negative shocks propagate through supply chains, affecting performance of firms linked through supply chains. Because the corporate value of publicly listed firms is evaluated in the stock market,[24, 25, 26] a shock to a firm is expected to affect not only its own stock price but also the stock prices of its suppliers and client firms. For example, when Boeing 737 MAX crashed in 2019 and 2020, the stock price of Boeing, its supplier, such as GE and Allegheny Technologies, and its clients, such as American Airlines and Southwest Airlines, dropped sharply.[27, 28, 29, 30, 31]. Also, a misconduct by Daihatsu Motor Co., a Japanese automobile manufacturer affiliated with Toyota, in 2023 negatively affected the stock price of Daihatsu, Toyota, and their suppliers and clients, such as Aisin.[32, 33]
@@
These examples show that the shock can propagate through supply chains and affect stock prices of firms linked through supply chains. Although relatively a lot of previous studies have examined the direct impact of shocks to firms on their own stock prices[34, 35, 36], there are not so many studies that investigated the diffusion of shocks to other firms’ stock prices through supply chains. Kim and Wagner [37] analyzed 315 cases reported by the Wall Street Journal regarding supply chain risks, which also included mentions of 69 suppliers and 246 clients. Pandit et al. utilized a larger dataset of quarterly earnings announcements, which differ from the news reports considered in our study, and analyzed their propagation across 1,518 suppliers and 4,010 clients in COMPUSTAT[38]. In addition, similar studies focusing on China are discussed [39, 40]. However, to the best of our knowledge, no study has examined large-scale data spanning more than a decade and covering tens of thousands of firms.
@@
To fill this gap, this study examines this possible diffusion, using two large firm-level samples, one for firms across the world and the other for Japanese firms. Specifically, we test whether the level of positive and negative sentiment about a firm in a news article affects the change rate of the stock price of the firm and its suppliers and clients before and after the disclosure of the news.
We contribute to the literature by utilizing large-scale data in many respects. First, our firm-level data covers most publicly listed firms in the world and in Japan and identify detailed supply chains among them using financial statements, news articles, websites, and firm-level surveys. Second, our data on news articles are taken from the NewsArchive of Thomson Reuters and cover more than 20 million articles. Third, we employ daily stock prices of all listed firms to examine both short- and long-run effects of positive news.
@@
Although we have mainly focused on how the shock affects stock prices through supply chains, it is evident that this study also relates to business applications and researches that predict changes of stock prices based on the news. In this field, neural networks are predominantly used. For example, hedge funds, such as Bridgewater Associates utilizes the neural networks for predictions[41]. In academia, the transformer[42] is widely used[43, 44]. In addition to the transformer, there is a long history of the usage of neural networks to predict stock prices, such as long short-term memory[45, 46, 47, 48, 49, 50, 51, 52], convolutional neural networks[53], or other deep learning methods[54, 55, 56, 57]. We use sentiment analysis to determine the positiveness and negativeness of each news article[58, 59] and for this purpose, specifically, Bidirectional Encoder Representations from Transformers (BERT)[60] has garnered the most acclaim[35]. We take advantage of FinBERT[61], a fine-tuned model of BERT particularly for financial information.
@@ 2 Data
This study utilizes four major data sources. First, our information on news relies on news articles written in English included in the NewsArchive provided by Thomson Reuters for the period 2003-2016. The information in each news article includes the time, headline, main text, Refinitiv Identification Code (a ticker-like code) of firms mentioned in the article, language, etc. As we will explain in detail later, our sentiment analysis, which determines the degree of positivity or negativity of each news article about the firms mentioned, utilizes the main text. The average number of words in an article is 243, whereas its median is 136 and maximum is 3,911. Although the total number of news articles in the NewsArchive for the period examined is 20,803,561, we focus on 3,447,425 that mention any firm publicly listed in 105 stock markets in the world. Among these news articles, 90% mention three or fewer firms, while the maximum number of firms mentioned in a single article is 33. (Figure 1). The number of news articles that mention a particular firm during the period 2003-2016 is quite skewed. 55.6% of firms were never mentioned, while 0.1% were mentioned more than 10,000 times (Figure 2).
@@
Second, we utilize daily stock prices of currently listed firms in stock markets in the world taken from the Eikon database provided by Refinitiv, one of the largest providers of financial markets data and infrastructure, using its Application Programming Interface (API). We drop firms listed in the past but not currently from our sample, because their stock prices are not available using the API. We particularly define the daily stock price of each firm as its closing price. In addition to stock prices of each firm, we utilize the Refinitiv index for each stock market that indicates the average of stock prices in the market to control for the market-specific time trend. The Refinitiv index is available for 105 major markets that cover 92% of listed firms worldwide.
@@
Third, to identify global supply chains among publicly listed firms in the world, we rely on data from FactSet that cover a number of firms across the world for the period 2003–2016 obtained as of January 18, 2023. The FactSet data include major clients of each firm reported in its financial statement. The US, Japanese, and many other governments require every listed firm to disclose its major clients with which the sales are more than 10% of its total sales, at least in the period examined in this study [62]. Therefore, based on published financial statements and, in addition, supplemental information from news articles and web sites, FactSet identifies supply chains of listed and non-listed firms with their major partners. Table 1 shows the number of all firms (for reference) and listed firms (our sample) in the FactSet data and their supply-chain links by year. The number of firms covered in the FactSet data increased rapidly over time, because the dataset has been constructed recently and thus ignores firms in the distant past. In a later section, the recorded supply chain links at the time are considered to estimate the impact of the news on supply chains.
@@
Finally, to identify domestic supply chains within Japan, we employ data collected by Tokyo Shoko Research (TSR) for the period 2008–2016, particularly its Company Linkage Database that covers supply chains of more than one million firms in Japan. Supply chains are captured by annual firm-level surveys by TSR that request each firm to report up to 23 suppliers and client firms of each firm. Apparently, many firms are linked with more than 23 suppliers and clients, and thus their links cannot be fully identified by their own responses. However, these links can be mostly identified by responses by their suppliers and clients. Since those firms are relatively large, their partners report them as important suppliers or clients. The number of firms and their supply-chain links and the corresponding number for listed firms from 2008 to 2016 are presented in Table 2.
@@ 3 Methods
Our sentiment analysis to determine the degree of positiveness and negativeness of the information about the listed firms provided in each news article utilizes a natural language processing (NLP) model called FinBERT[61]. FinBERT is a variant of the Bidirectional Encoder Representations from Transformers (BERT) developed by Devlin et al. (2018)[60]. Specifically, FinBERT uses Reuters’ TRC2-financial that consists of 1.8 million news articles between 2008 and 2010 and fine-tunes BERT for financial sentiment classification using Financial Phrasebank that consists of 4,845 English sentences taken from financial news in the LexisNexis database. From each news article, FinBERT generates softmax outputs for three labels, i.e., positive, neutral, and negative, which indicate the weights of the three about the firms mentioned in the article. Examples are shown in Table 3. Before conducting the sentiment analysis using FinBERT, we clean news articles by deleting URLs, line breaks, ISINs, and fixed phrases at the end. The distribution of the probabilities of positive, neutral, and negative sentiments generated by FinBERT from news articles about firms across the world and Japanese firms are provided in Panels (A) and (B) of Figure 3, respectively. Both panels indicate that the probability of neutral sentiment is high for many news articles followed by that of negative sentiment while news articles with a large probability of positive sentiment are fewer.
@@ 3.1 Regression analysis
Direct effect
To estimate the effect of positive and negative sentiment in news articles about particular firms on their own stock prices, we apply the probabilities of positive and negative sentiment constructed above to regression analysis. Specifically, we focus on firms mentioned by any news article and examine whether the change rate of their stock prices in a time window are affected by the level of positiveness or negativeness of the article after the article is disclosed, compared with its change rate before the disclosure. Suppose that firm listed in stock market is mentioned by news article on day . Setting the time window at , such as 1, 5, and 30 days, we define the average daily percentage change of the stock price of firm before and after news article disclosed on day in time window , or and , respectively, by
@@
where T is either pre or post, and and are indicator variables that take a value of one if represents the percentage change in the stock price of firm before or after news article is disclosed on day , respectively. represents the level of either positive or negative sentiment in news article generated by FinBERT (Data section). is the percentage change of the Refinitiv index for each stock market that controls for the overall time trend of stock prices in the market, and is sector dummies. Sectors are defined by the North American Industry Classification System at the sub-sector level for the sample of firms across the world and by the Japan Standard Industrial Classification at the two-digit level for the sample of Japanese firms.
Accordingly, indicates the average effect of the positiveness or negativeness of a news article on the average change rate of the stock price of the firm mentioned in the article after it is disclosed in the time window of days, controlling for the overall trend in the market and sector-specific unobservable factors. By contrast, indicates any prior “effect” of the news article on the change rate of the stock price, including the possible effect of the information in the news article that diffuses in the market not through the news but through unofficial information channels before the disclosure of the news. Such unofficial information channels include suggestions by financial companies to their clients, posts in Social Networking Services (SNS) by influencers, and word of mouth among investors. In addition, may include the endogeneity bias reflecting the possibility that the growth rate of the stock price of a firm that is positively mentioned by any news article is intrinsically higher than that of a firm negatively mentioned. Therefore, the net effect of the disclosure of the information by the news article on the change rate of the stock price in time window should be given by [66].
We particularly estimate equation (3) for 1, 2, 3, 4, 5, 30, 180, and 365 to examine both the short- and long-run effect of positive news articles on stock prices. We hypothesize that a positive (negative) news article would accelerate (deteriorate) the change rate of the stock price of the firm mentioned in the article in the short-run, while the effect would become smaller and converge to 0 in the longer run.
@@
Indirect effect through supply chains
We further examine the effect of positive or negative sentiment about firms in news articles on their suppliers and clients, using similar frameworks. Specifically, when we focus on the upstream effect of news articles about firms on their suppliers, in equation (3) is replaced with the probability of positive or negative sentiment in the news article that mentions not firm but any of ’s client firms. Alternatively, when we examine the downstream effect of clients’ news on their clients, is replaced with the positive or negative index of any of the suppliers of firm . When a firm mentioned by a news article has more than one supplier or client, as shown in Tables 1 and 2, we include all the suppliers and clients in our sample. In this analysis, we compare stock prices of suppliers or clients of firms mentioned positively in news articles with those of suppliers or clients of firms mentioned negatively, as in our analysis on the direct effect on own stock prices.
We hypothesize that the effect of news articles on stock prices diffuses through supply chains. Precisely, a positive news article about a client firm would accelerate the growth in the stock price of its suppliers, in addition to its own stock price, because the market predicts an increase in demand for products of the suppliers associated with higher performance in the client firm. Also, positive news about a supplier would promote growth of its clients because of possible productivity growth of the supplier. As in the case of the direct effect, the indirect effect on suppliers and clients would be smaller in the long run than in the short run.
4 Results and Discussion
Direct effect of news articles on listed firms across the world
We start with estimating how the positiveness of news articles about firms affect percentage changes in their stock prices before and after the disclosure of news, applying equation (3) to the sample of publicly listed firms across the world mentioned by any news article. Results are presented in Table 4 and Figure 5.

### v4

# Stock Price Responses to Firm-Level News in Supply Chain Networks

@@ Abstract
This study examines how positive and negative news about firms are associated with stock prices and whether these associations extend to suppliers and clients linked via supply chain relationships, using large samples of publicly listed firms worldwide and in Japan. News sentiment is measured using FinBERT, a natural language processing model fine-tuned for financial text, and supply chain links are identified from financial statements for global firms and from large-scale firm-level surveys for Japanese firms. We find that stock prices exhibit systematic associations with positive and negative news even before public disclosure. These associations are also observed for suppliers and clients before and after disclosure. In general, post-disclosure associations are larger than pre-disclosure associations, with the difference concentrated around the time of public news disclosure relative to the pre-disclosure period. However, for Japanese firms, the post-disclosure associations for suppliers and clients are smaller than the pre-disclosure associations, in contrast to the pattern observed for firms outside Japan.
1 Introduction
@@
Firms interact with and influence each other through various types of networks. Major firm networks that are found to play a significant role in firms’ behaviors and performance in the literature include financial networks through ownership [1, 23, 27] and knowledge networks through research collaboration [30, 29, 10, 72, 34]. Another type of network recently paid attention to in the academic literature and business and policy fields is supply chains formed through transactions of materials, parts, and components [15, 32]. One reason for the growing attention is that supply chains have expanded globally, linking firms in the world within a small number of steps with each other [6].
@@
Global supply chains have rapidly expanded because firms interconnected through supply chains can greatly benefit from the network. Most notably, supply chains can empower firms by enhancing the efficiency of production processes and thus optimizing firms’ economic performance [11, 5]. For instance, by procuring and manufacturing different goods in various locations based on each location’s comparative advantages and cost-effectiveness, firms can streamline their operations as a network, ultimately leading to reductions in production costs and improvements in productivity [18]. Moreover, it is often evident that productivity spills over through supply chains from upstream suppliers to downstream client firms because of high-quality materials, parts, components, and services and from downstream to upstream firms because of learning of knowledge and technology [55, 72, 38].
@@
Conversely, supply chains can be a channel of negative shocks that aggravate firm performance, particularly once supply chains are disrupted due to unexpected events, such as natural disasters and geopolitical instability [9, 35, 7]. Such negative shocks as examined in the literature include the Great East Japan Earthquake in 2011 [49], the massive flood in Thailand in 2011 [31], the COVID-19 pandemic from 2020 to 2022 [17], and the Russia-Ukraine war that started in 2022 [54].
@@
Accordingly, prior studies show that positive or negative shocks propagate through supply chains, affecting the performance of firms linked through supply chains [7]. Because the corporate value of publicly listed firms is evaluated in the stock market [52, 71, 12], a shock to a firm is expected to affect not only its own stock price but also the stock prices of its suppliers and client firms. For example, following the Boeing 737 MAX crashes in 2019 and 2020, stock prices of Boeing, some of its suppliers such as GE and Allegheny Technologies, and some of its clients such as American Airlines and Southwest Airlines declined sharply [13, 14, 22, 48, 62]. Also, a misconduct by Daihatsu Motor Co., a Japanese automobile manufacturer affiliated with Toyota, in 2023 was accompanied by stock price declines for Daihatsu, Toyota, and related firms including suppliers and clients such as Aisin [64, 63].
@@
These examples show that the shock can propagate through supply chains and affect stock prices of firms linked through supply chains. Although relatively a lot of previous studies have examined the direct impact of shocks to firms on their own stock prices [44, 69, 68], there are not so many studies that have investigated the diffusion of shocks to other firms’ stock prices through supply chains. Cohen and Frazzini [16] show that information about a firm’s customers predicts future stock returns of its suppliers, providing early evidence that stock price reactions propagate through supply chain relationships. More recently, Le Tran and Coqueret [42] show that firm-level news shocks, particularly ESG-related news, affect not only the focal firm’s stock returns but also the returns of its suppliers and clients, providing direct evidence of news-driven spillovers along supply chain relationships. Complementing this evidence, Agarwal et al. [2] demonstrate that investor attention, measured through online co-search behavior, facilitates information diffusion across supply chain partners and contributes to predictable stock return comovements. Kim and Wagner [40] analyzed 315 cases reported by the Wall Street Journal regarding supply chain risks, which also included mentions of 69 suppliers and 246 clients. Pandit et al. utilized a larger dataset of quarterly earnings announcements, which differ from the news reports considered in our study, and analyze their propagation across 1,518 suppliers and 4,010 clients in COMPUSTAT [57]. In addition, similar studies focusing on China are discussed [45, 46]. However, to the best of our knowledge, no study has examined large-scale data spanning more than a decade and covering tens of thousands of firms.
@@
To fill this gap, this study examines whether news sentiment is associated with stock price changes of firms and of their suppliers and clients in the pre- and post-disclosure periods, using two large firm-level samples, one for firms across the world and the other for Japanese firms. Specifically, we test whether the level of positive and negative sentiment about a firm in a news article affects the change rate of the stock price of the firm and its suppliers and clients before and after the disclosure of the news.
@@
Beyond supply-chain-specific mechanisms, a growing literature examines how the content of news itself is quantified and linked to stock market reactions. Early studies quantified news content using dictionary-based and word-frequency approaches to capture investor sentiment and market reactions. For example, Tetlock [70] construct a sentiment measure based on word categories in financial news and show that media pessimism predicts short-run market returns and trading volume. Relatedly, Mizuno et al. [53] quantify the novelty and topicality of business news using word-level similarity measures and demonstrate that such text-based characteristics are strongly associated with subsequent stock market activity. In this field, neural networks are predominantly used. For example, hedge funds, such as Bridgewater Associates, utilize neural networks for predictions [8]. In academia, the transformer [73] is widely used [67, 25]. In addition to the transformer, there is a long history of the usage of neural networks to predict stock prices, such as long short-term memory [37, 41, 20, 21, 39, 43, 65, 26], convolutional neural networks [33], or other deep learning methods [47, 58, 66, 56]. We use sentiment analysis to determine the positiveness and negativeness of each news article [51, 28] and for this purpose, specifically, Bidirectional Encoder Representations from Transformers (BERT) [19] have garnered the most acclaim [69]. We take advantage of FinBERT [4], a fine-tuned model of BERT particularly for financial information.
To advance this line of research from market- or firm-level settings to supply-chain networks and cross-country comparisons, it is essential to combine large-scale news data with comprehensive firm-level linkages and stock price information. We contribute to the literature by utilizing large-scale data in many respects. First, our firm-level data covers most publicly listed firms in the world and in Japan and identifies detailed supply chains among them using financial statements, news articles, websites, and firm-level surveys. Second, our data on news articles are taken from the News Archive of Thomson Reuters and cover more than 20 million articles. Third, we employ daily stock prices of all listed firms to examine both short- and long-run effects of positive and negative news.
@@ 2 Data
2.1 News article
Our information on news relies on news articles written in English included in the News Archive provided by Thomson Reuters for the period 2003–2016. In the News Archive, multiple news articles may be filed for a single event. Typically, the first article is an alert, which is a short message containing only essential facts. This is followed by a newsbreak, usually released within 5–20 minutes after the alert. A newsbreak consists of a headline, which is similar to that of the alert, and two to four paragraphs of body text. Subsequent updates may follow the newsbreak after an additional 20–30 minutes, providing further information.
In this study, we use the newsbreak associated with each event. Although we are interested in the timing of public disclosure, alerts often do not include sufficient details about the related firms, whereas newsbreaks do. The time difference between alerts and newsbreaks is typically less than 20 minutes, which is negligible given that our analysis is conducted at the daily frequency. Accordingly, the release time of the newsbreak is aligned with the closest market closing time of the focal stock market, and the corresponding trading day is defined as day zero in the subsequent analyses. The newsbreak includes the time, headline, main text, Refinitiv Identification Codes (RICs) of one or more listed firms mentioned in the article, the language, and other metadata.
@@ 2.1 News article
As we will explain in detail later, our sentiment analysis, which determines the degree of positivity or negativity of each news article about the firms mentioned, utilizes the main text. The average number of words in an article is 243, whereas its median is 136 and its maximum is 3,911. Although the total number of news articles in the News Archive for the period examined is 20,803,561, we focus on 3,447,425 that mention any firm publicly listed in 105 stock markets in the world. Among these news articles, 90% mention three or fewer firms, while the maximum number of firms mentioned in a single article is 33 (Figure 1). The number of news articles that mention a particular firm during the period 2003-2016 is quite skewed. 55.6% of firms were never mentioned, while 0.1% were mentioned more than 10,000 times (Figure 2).
Stock price
@@
We utilize daily stock prices of currently listed firms in global stock markets, obtained from the Eikon database provided by Refinitiv, one of the largest providers of financial markets data and infrastructure, through its Application Programming Interface (API). Firms that were listed in the past but are no longer listed are excluded from the sample, as their stock price data are not available through the API. We define the daily stock price of each firm as its closing price. Accordingly, the stock price data include Refinitiv Identification Codes (RICs), dates, and closing prices.
In addition to stock prices of each firm, we utilize the Refinitiv index for each stock market that indicates the average of stock prices in the market to control for the market-specific time trend. The Refinitiv index is available for 105 major markets that cover 92% of listed firms worldwide.
Supply chain
@@
To identify global supply chains among publicly listed firms in the world, we rely on data from FactSet that cover a number of firms across the world for the period 2003–2016 obtained as of January 18, 2023. The FactSet data include major clients of each firm reported in its financial statements. The US, Japanese, and many other governments require every listed firm to disclose its major clients with which the sales are more than 10% of its total sales, at least in the period examined in this study [24]. Therefore, based on published financial statements and, in addition, supplemental information from news articles and websites, FactSet identifies supply chains of listed and non-listed firms with their major partners. Table 1 shows the number of all firms (for reference) and listed firms (our sample) in the FactSet data and their supply-chain links by year. The number of firms covered in the FactSet data increased rapidly over time, because the dataset has been constructed recently and thus ignores firms in the distant past. In a later section, the recorded supply chain links at the time are used to examine news-related stock price changes along supply chain relationships.
The FactSet supply chain database does not cover the universe of firms worldwide. Instead, it identifies major customer–supplier relationships disclosed in financial statements and related sources, which are typically limited to economically significant partners. As a result, its coverage is intentionally limited, especially in earlier years.
FactSet uses its own identifier system, referred to as the FactSet Security Identifier. We convert these identifiers to Refinitiv Identification Codes (RICs) in order to link the FactSet data with news and stock price data.
@@
To identify domestic supply chains within Japan, we employ data collected by Tokyo Shoko Research (TSR) for the period 2008–2016, particularly its Company Linkage Database that covers supply chains of more than one million firms in Japan. Supply chains are captured by annual firm-level surveys by TSR that request each firm to report up to 23 suppliers and client firms for each surveyed firm. Apparently, many firms are linked with more than 23 suppliers and clients, and thus their links cannot be fully identified by their own responses. However, these links can be mostly identified by responses by their suppliers and clients. Since those firms are relatively large, their partners report them as important suppliers or clients. Therefore, although the numbers of suppliers and clients are limited to 23 in the reports, large firms can have many partners. Accordingly, the maximum number of suppliers and clients is 12,729 and 12,016, respectively, in 2016, providing broad coverage of supply chains of Japanese firms.
The TSR database provides exceptionally broad coverage of firm-level relationships, including both listed and unlisted firms. In this study, however, we restrict our analysis to relationships involving listed firms and links between such firms, in order to ensure consistency with market-based information. Listed firms in the TSR database are identified by corporate IDs and converted into Refinitiv Identification Codes (RICs). The number of firms and their supply-chain links and the corresponding number for listed firms from 2008 to 2016 are presented in Table 2.
@@ 3 Methods
Our sentiment analysis to determine the degree of positiveness and negativeness of the information about the listed firms provided in each news article utilizes a natural language processing (NLP) model called FinBERT [4]. FinBERT is a variant of the Bidirectional Encoder Representations from Transformers (BERT) developed by Devlin et al. (2018) [19]. Specifically, FinBERT uses Reuters’ TRC2-financial, which consists of 1.8 million news articles between 2008 and 2010 and fine-tunes BERT for financial sentiment classification using Financial Phrasebank, which consists of 4,845 English sentences taken from financial news in the LexisNexis database. From each news article, FinBERT generates softmax outputs for three labels, i.e., positive, neutral, and negative, which indicate the weights of the three sentiments about the firms mentioned in the article. Examples are shown in Table 3. Before conducting the sentiment analysis using FinBERT, we clean news articles by deleting URLs, line breaks, ISINs, and fixed phrases at the end. The distribution of the probabilities of positive, neutral, and negative sentiments generated by FinBERT from news articles about firms across the world and Japanese firms are provided in Panels (A) and (B) of Figure 3, respectively. Both panels indicate that the probability of neutral sentiment is high for many news articles followed by that of negative sentiment while news articles with a large probability of positive sentiment are fewer.
From the three sentiment probabilities generated by FinBERT (positive, neutral, and negative), we construct sentiment measures by excluding the neutral component and normalizing the remaining positive and negative probabilities. Specifically, the positive sentiment index is defined as
and the negative sentiment index as
where and denote the probabilities of positive and negative sentiment, respectively, for news article mentioning firm on day . This normalization focuses the analysis on the relative balance between positive and negative information and avoids ambiguity associated with neutral sentiment.
@@ 3.1 Regression analysis
Own-firm associations
To examine the association between positive and negative sentiment in news articles about particular firms and their own stock prices, we apply the probabilities of positive and negative sentiment constructed above to regression analysis. Specifically, we focus on firms mentioned by any news article and examine whether the change rate of their stock prices in a time window is associated with the level of positiveness or negativeness of the article after the article is disclosed, compared with its change rate before the disclosure. Suppose that firm listed in stock market is mentioned by news article on day . Setting the time window at , such as 1, 5, and 30 days, we define the average daily percentage change of the stock price of firm before and after news article was disclosed on day in time window , or and , respectively, by
@@
where T is either pre or post, and and are indicator variables that take a value of one if represents the percentage change in the stock price of firm before or after news article is disclosed on day , respectively. represents the level of positive or negative sentiment in news article generated by FinBERT and corresponds to the sentiment indices and defined in the Methods section. is the percentage change of the Refinitiv index for each stock market that controls for the overall time trend of stock prices in the market, and is sector dummies. Sectors are defined by the North American Industry Classification System at the sub-sector level for the sample of firms across the world and by the Japan Standard Industrial Classification at the two-digit level for the sample of Japanese firms.
Accordingly, indicates the average association between the positiveness or negativeness of a news article and the average change rate of the stock price of the firm mentioned in the article after it is disclosed in the time window of days, controlling for the overall trend in the market and sector-specific unobservable factors. By contrast, captures any prior association between the news article and the change rate of the stock price before disclosure. This term may reflect a combination of factors that are not separately identified in this analysis, including selection into news coverage and pre-existing price dynamics. In addition, may include the endogeneity bias reflecting the possibility that the growth rate of the stock price of a firm that is positively mentioned by any news article is intrinsically higher than that of a firm negatively mentioned. Accordingly, the difference between the post- and pre-news coefficients, , summarizes changes in stock price growth around the time of news disclosure [36], relative to the pre-disclosure period. This difference compares post-disclosure price movements with those observed before disclosure, where pre-news movements may reflect the incorporation of information prior to public disclosure.
We particularly estimate equation (5) for 1, 2, 3, 4, 5, 30, 180, and 365 to examine both the short- and long-run associations between positive news articles and stock prices. We hypothesize that a positive (negative) news article is associated with higher (lower) stock price changes in the short-run, with weaker associations in the longer run.
@@
Associations along supply chains
We further examine the association between positive or negative sentiment about firms in news articles and stock price changes of their suppliers and clients, using similar frameworks. Specifically, when we focus on upstream associations between news articles about firms and their suppliers, the variable in equation (5) is replaced with the positive or negative sentiment indices from news articles that mention any of firm ’s client firms. Alternatively, when we examine the downstream associations between news articles about firms and their clients, is replaced with the positive or negative indices of any of the suppliers of firm . When a firm mentioned by a news article has more than one supplier or client, as shown in Tables 1 and 2, we include all the suppliers and clients in our sample.
We hypothesize that news sentiment about a firm is associated with stock price changes of its suppliers and clients. More precisely, news sentiment about a client firm is associated with stock price changes of its suppliers, in addition to those of the client firm itself, while news sentiment about a supplier firm is associated with stock price changes of its clients. These associations may reflect, for example, investors’ expectations about changes in demand faced by suppliers or about the performance of downstream firms, although such mechanisms are not directly tested in this study. As in the case of own-firm associations, these associations are expected to be stronger over shorter horizons and weaker over longer horizons.
4 Results
Own-firm stock price associations for listed firms across the world
We start by examining how the positiveness of news articles about firms is associated with percentage changes in their stock prices before and after the disclosure of news, applying equation (5) to the sample of publicly listed firms across the world mentioned by any news article. Results are presented in Table 4 and Figure 5. Then, we show the results for the negativeness.